# Identifying multi-compartment Hodgkin-Huxley models with high-density extracellular voltage recordings

### v1

@@ Abstract
1 Department of Statistics and the Wu Tsai Neurosciences Institute, Stanford University, Stanford, CA, USA
2 Machine Learning in Science, University of Tübingen, Germany
3 Tübingen AI Center, Tübingen, Germany
4 Department of Empirical Inference, Max Planck Institute for Intelligent Systems, Tübingen, Germany
Correspondence: ictanoh@stanford.edu and scott.linderman@stanford.edu
@@ 1 Introduction
Biological neurons are equipped with a wide range of biophysical mechanisms that contribute to information processing (Poirazi et al., 2003a; Spruston, 2008; Gidon et al., 2020) and to the brain’s robustness to environmental perturbations (O’Leary et al., 2013; Haddad and Marder, 2018). Multi-compartment Hodgkin-Huxley models provide a principled framework for studying biophysical mechanisms of neurons (Hodgkin and Huxley, 1952; Carnevale and Hines, 2006). Multi-compartment models treat neurons as connected electrical compartments, allowing simulation of how voltage signals propagate throughout a cell as ion channels open and close. A central limitation for building multi-compartment models—and for understanding biophysical mechanisms in neurons—is the lack of experimental techniques to precisely measure the biophysical parameters of these models. As such, these parameters must be inferred from recordings of the activity of neurons (Huys et al., 2006; Mensi et al., 2012; Almog and Korngreen, 2014; Gonçalves et al., 2020). Typically, these recordings are made from within a neuron using intracellular measurements such as patch-clamp electrophysiology, which enables precise recordings of voltage dynamics in the soma (Sakmann and Neher, 1984; Hamill et al., 1981) and, for large neurons, in dendrites (Stuart and Sakmann, 1994). However, intracellular methods are technically challenging, and only provide voltage signals from a very small number of locations within a neuron (typically, only from the soma) (Margrie et al., 2002).
Recent advances in neural recording technologies have opened new possibilities for studying neuron biophysics in vivo. High-density silicon probes such as the Neuropixel Ultra (Ye et al., 2024) can record extracellular voltage traces across hundreds of closely spaced electrodes. These signals, which reflect the transmembrane currents generated by nearby neurons, are traditionally used for spike detection and sorting—identifying when and where neurons fire spikes (Pachitariu et al., 2016; Rossant et al., 2016).
@@
While effective for tracking population activity, this approach treats spikes as discrete events and ignores the exquisite biophysical details of neurons (Einevoll et al., 2012).
In this work, we revisit the extracellular signal, not just as a source of discrete spike times, but as a rich, spatially structured observation from which we can infer biophysical properties of individual neurons. This is made possible by three key developments: (1) high-density probes and 2D arrays that allow us to sample extracellular voltage at many sites around a cell to triangulate its position and morphology; (2) differentiable simulators like Jaxley (Deistler et al., 2024) that enable fast, hardware-accelerated simulation of complex biophysical models and gradient-based parameter optimization; and (3) scalable state-space inference frameworks such as Dynamax (Linderman et al., 2025), which allow efficient state inference and parameter estimation in high-dimensional, nonlinear dynamical systems.
We leverage these developments to provide a statistical approach for inferring the parameters of multi-compartment biophysical models from extracellular observations: We extend classical multi-compartment Hodgkin–Huxley models to include a biophysical model of extracellular potentials. We then use an Extended Kalman Filter (EKF) (Särkkä and Svensson, 2023) to infer latent neuronal states—such as membrane voltages and gating variables—from observed extracellular signals and we improve the computational efficiency with a diagonal approximation to the filtering covariance. To estimate model parameters, including ion channel densities and the neuron’s spatial position relative to the probe, we maximize the marginal likelihood using gradient-based optimization.
@@ 2.1 Modeling Action Potentials with Hodgkin–Huxley Dynamics
where $I_{\text{ext}}^{(i)}$ is the external current applied to compartment $i$, and $g^{(i,j)}$ is the axial conductance coupling compartment $i$ to its neighbor $j$. The total transmembrane current in compartment $i$ is denoted by ${I_{m}^{(i)}=\sum_{c\in\mathcal{C}}g_{c}^{(i)}a_{c}(\lambda_{c}^{(i)})(V^{(i)}-E_{c})}$. For instance, in the original single compartment Hodgkin–Huxley model with sodium, potassium, and leak channels, the total current is,
$I_{m}=g_{\text{Na}}\,m^{3}h(V-E_{\text{Na}})+{g}_{\text{K}}\,n^{4}(V-E_{\text{K}})+{g}_{\text{L}}(V-E_{\text{L}}).$ (2)
Here, for example, the states of the sodium channel are $\lambda_{\mathrm{Na}}=(m,h)\in[0,1]^{2}$ and the activation function is $a_{\mathrm{Na}}(\lambda_{\mathrm{Na}})=m^{3}h$. The gating variables evolve according to first-order kinetics $\frac{\mathrm{d}\lambda_{c}}{\mathrm{d}t}=\alpha_{c}(V)(1-\lambda_{c})-\beta_{c}(V)\lambda_{c},$ where $\alpha_{c}(V)$ and $\beta_{c}(V)$ are voltage-dependent rate functions that control the opening and closing transitions of the ion channel. Specific forms for the $\alpha$ and $\beta$ functions are available in Dayan and Abbott (2005), e.g.
@@ 2.2 Modeling Extracellular Voltage Measurements
Extracellular voltages reflect the collective activity of nearby neurons. While the physics of extracellular recordings are well understood (Gold et al., 2006; Lindén et al., 2014), our work is, to our knowledge, the first to leverage this framework for fitting multi-compartment Hodgkin–Huxley models to extracellular measurements. Let $\Phi$ denote the extracellular potential measured by a fixed site on a recording probe. We treat each compartment as a point source of transmembrane current located at its center and denote $r_{i}$ the distance from the recording site to compartment $i$ of a neuron. It follows from Coulomb’s law that the measured extracellular voltage is,
@@
In words, the observed potential $\Phi$ is proportional to the sum of transmembrane currents from nearby compartments, weighted by the inverse of the distance to the compartment.
@@ 3 State Inference and Parameter Estimation
We aim to infer the membrane potential and channel states and to estimate the biophysical and structural parameters of a neuron from noisy extracellular voltage traces. We can achieve both goals by casting the multi-compartment Hodgkin-Huxley model as a state space model (SSM) with nonlinear dynamics. However, the nature and scale of these models present unique challenges, which we address below.
@@ 3.1 State Space Model Formulation
A state space model (SSM) is a latent variable model with latent states $\mathbf{z}_{1:T}$, observations $\mathbf{y}_{1:T}$, and parameters $\theta$. In a Markovian SSM, latent states evolve according to a dynamics distribution, $p(\mathbf{z}_{t}~{}|~{}\mathbf{z}_{t-1};\theta)$, and they give rise to observations via an emission distribution $p(\mathbf{y}_{t}~{}|~{}\mathbf{z}_{t};\theta)$. There are a host of well-established techniques for state inference and parameter estimation in such models (Särkkä and Svensson, 2023; Linderman et al., 2025), depending on the nature of the dynamics and emission distributions.
In our case, the model parameters, $\theta$, may include the maximum conductances of various ion channels as well as parameters governing the neuron’s spatial configuration relative to the recording probe. The latent states consist of the voltages and channel states for each compartment, $\mathbf{z}_{t}=[V^{(1)},\dots,V^{(N)},\{\lambda_{c}^{(1)}\}_{c\in\mathcal{C}_{1}},\dots,\{\lambda_{c}^{(N)}\}_{c\in\mathcal{C}_{N}}]$. To model the state dynamics, we use an Euler discretization of the multi-compartment Hodgkin-Huxley differential equations (see eq. 1),
$\mathbf{z}_{t+1}$ $\sim\mathcal{N}(f_{\theta}(\mathbf{z}_{t},t),\,dt\,\Sigma_{\text{dyn}}),$ (4)
where $f_{\theta}(\mathbf{z}_{t},t)$ is the deterministic update function derived from the ODEs above, and $\Sigma_{\text{dyn}}$ is a diagonal covariance matrix with entries from $\{\sigma^{2}_{v},~{}\sigma^{2}_{\text{state}}\}$. This Gaussian noise is intended to account for model uncertainty, such as missing ion channels or misspecified channel dynamics, which we will consider in the results section. Since the gating variables are constrained to the interval $[0,1]$, they are transformed using the logit function. For simplicity, we also refer to the transformed vector as $\mathbf{z}_{t}$.
[…]
$\mathbf{y}_{t}\sim\mathcal{N}(h_{\theta}(\mathbf{z}_{t}),\sigma^{2}_{\text{obs}}I),$ (5)
where $h_{\theta}(\mathbf{z}_{t})$ is the predicted extracellular voltage from eq. 3 and $\sigma^{2}_{\text{obs}}$ is the observation noise variance. Since $h_{\theta}(\mathbf{z}_{t})$ only depends on $\mathbf{z}_{t}$, the emission model is conditionally independent of previous states given the current latent state.
@@ 3.2 Scaling the Extended Kalman Filter to Large Biophysical Models
Since the dynamics and emission models are nonlinear, there is no closed-form solution for the filtering distribution $p_{\theta}(\mathbf{z}_{1:t}~{}|~{}\mathbf{y}_{1:t})$. To address this challenge, we use the extended Kalman filter (EKF), which linearizes the dynamics and emission functions via first-order Taylor expansions (Särkkä and Svensson, 2023). This linear approximation enables recursive estimation of the latent states, and it also provides an estimate of the marginal likelihood $p_{\theta}(\mathbf{y}_{1:T})$ for parameter estimation.
@@
Biologically realistic neuron models often contain dozens or even hundreds of compartments, leading to latent state dimensions in the hundreds or thousands. The standard EKF maintains an estimate of the full state covariance matrix, and updating the filtering distribution requires solving a linear system with this matrix. Since this computation scales cubically with the number of states, it becomes computationally infeasible to apply standard EKF formulations to full-scale neuronal models. To overcome this challenge, we develop a scalable EKF variant for biophysical neuron model fitting. Our approach leverages a diagonal approximation of the posterior covariance, inspired by Chang et al. (2022), as well as the special structure of multi-compartment Hodgkin-Huxley dynamics.
We approximate the EKF filtering distribution $p_{t}(x)=\mathcal{N}(x;\mu_{t},\mathbf{\Sigma}_{t})$ with a diagonal Gaussian distribution $q_{t}(x)=\mathcal{N}(x;\tilde{\mu}_{t},\tilde{\mathbf{\Sigma}}_{t})$ that minimizes the KL divergence $\text{KL}(q_{t}\,\|\,p_{t})$. This divergence minimization yields a closed-form solution: we match the marginal means and precisions (inverse covariances) of $p_{t}$ and $q_{t}$. The resulting diagonal covariance estimate is given by
$\tilde{\mathbf{\Sigma}}_{t}=\mathrm{diag}\left(\left[\left(\mathbf{\Sigma}^{ii}_{t|t-1}\right)^{-1}+\left(\mathbf{H}_{t}^{T}\mathbf{R}_{t}^{-1}\mathbf{H}_{t}\right)_{ii}\right]^{-1}\right),$ (6)
where $\mathbf{\Sigma}_{t|t-1}$ is the predicted covariance from the EKF prediction step, $\mathbf{R}_{t}=\sigma^{2}_{\text{obs}}I$ is the observation noise covariance, and $\mathbf{H}_{t}=\text{Jac}(h_{\theta})(\mu_{t~{}|~{}t-1})$ is the Jacobian of the emission function $h_{\theta}$ evaluated at the predicted state $\mu_{t~{}|~{}t-1}$.
Importantly, the diagonal of $\mathbf{H}_{t}^{T}\mathbf{R}_{t}^{-1}\mathbf{H}_{t}$ can be computed in linear time with respect to the number of states. The remaining challenge is to efficiently compute the diagonal of the predictive covariance, $\mathbf{\Sigma}_{t|t-1}$, given by
$\mathbf{\Sigma}_{t|t-1}=\mathbf{F}_{t}\mathbf{\Sigma}_{t-1}\mathbf{F}_{t}^{T}+\mathbf{Q}_{t},$ (7)
@@
where $\mathbf{F}_{t}=\text{Jac}(f_{\theta})(\mu_{t~{}|~{}t-1})$ is the Jacobian of the dynamics function $f_{\theta}$ evaluated at the predicted state $\mu_{t~{}|~{}t-1}$, and $\mathbf{Q}_{t}=dt\Sigma_{\text{dyn}}$ is the dynamics covariance matrix.
Chang et al. (2022) assumed $\mathbf{F}_{t}$ to be the identity matrix, simplifying the diagonal computation. In contrast, our dynamics are nonlinear, and $\mathbf{F}_{t}$ is nontrivial. However, assuming that $\mathbf{\Sigma}_{t-1}$ is diagonal, we can leverage the sparsity of $\mathbf{F}_{t}$ to compute the diagonal of $\mathbf{\Sigma}_{t~{}|~{}t-1}$ efficiently. In particular, while voltage updates may depend on all other states due to the implicit Euler solver used in Jaxley (Deistler et al., 2024), gating variables typically depend only on their own previous value and the voltage of their corresponding compartment. This localized structure implies that $\mathbf{F}_{t}$ is sparse and structured, allowing us to reduce the cost of computing the diagonal of $\mathbf{\Sigma}_{t~{}|~{}t-1}$ from quadratic to linear in the number of states (see Appendix F). Importantly, we do not compute the full dense Jacobian $\mathbf{F}_{t}$; instead, we use the sparsejac library (Schubert, 2023), which leverages graph coloring strategies to efficiently compute sparse Jacobians given a known sparsity pattern. This approach significantly reduces memory usage and computation time, making large-scale EKF updates tractable. With both diagonal components available in linear time, we can compute $\tilde{\mathbf{\Sigma}}_{t}$ efficiently, enabling scalable approximate EKF updates for large models.
Finally, to optimize model parameters, we perform gradient ascent on the marginal likelihood using the Adam optimizer (Kingma and Ba, 2014). We obtain an estimate for the marginal likelihood based on the EKF approximation,
$p(\mathbf{y}_{0:T})\approx\mathcal{N}(\mathbf{y}_{0};\mu_{0},\Sigma_{0})\prod_{t=1}^{T}\mathcal{N}\left(\mathbf{y}_{t};h_{\theta}(\mu_{t|t-1}),\mathbf{H}_{t}\mathbf{\Sigma}_{t|t-1}\mathbf{H}_{t}^{T}+\mathbf{R}_{t}\right),$ (8)
where $\mu_{0}$ and $\Sigma_{0}$ denote the initial state mean and covariance matrix respectively. Since we approximate $\mathbf{\Sigma}_{t~{}|~{}t-1}$ by its diagonal to construct $\tilde{\mathbf{\Sigma}}_{t}$, we also substitute this diagonal approximation into eq. 8 to estimate the marginal likelihood. Performing gradient ascent on the marginal likelihood requires to obtain the gradient through the biophysical model. To obtain this gradient, we use the Jaxley toolbox, a differentiable simulator of multi-compartment neuron models. All code used for inference, cell and voltage simulation, and the experiments presented in this paper is available at: https://github.com/ianctanoh/eap-fit-hh. Details of the computational resources used in this work are provided in Appendix A.

### v2

@@ 1 Introduction
Biological neurons are equipped with a wide range of biophysical mechanisms that contribute to information processing [poirazi2003pyramidal, spruston2008pyramidal, gidon2020dendritic] and to the brain’s robustness to environmental perturbations [o2013correlations, haddad2018circuit]. Multi-compartment Hodgkin-Huxley models provide a principled framework for studying biophysical mechanisms of neurons [hodgkin1952quantitative, carnevale2006neuron]. Multi-compartment models treat neurons as connected electrical compartments, allowing simulation of how voltage signals propagate throughout a cell as ion channels open and close. A central limitation for building multi-compartment models—and for understanding biophysical mechanisms in neurons—is the lack of experimental techniques to precisely measure the biophysical parameters of these models. As such, these parameters must be inferred from recordings of the activity of neurons [huys2006efficient, mensi2012parameter, almog2014quantitative, gonccalves2020training]. Typically, these recordings are made from within a neuron using intracellular measurements such as patch-clamp electrophysiology, which enables precise recordings of voltage dynamics in the soma [sakmann1984patch, hamill1981improved] and, for large neurons, in dendrites [stuart1994active]. However, intracellular methods are technically challenging, and only provide voltage signals from a very small number of locations within a neuron (typically, only from the soma) [margrie2002vivo].
Recent advances in neural recording technologies have opened new possibilities for studying neuron biophysics in vivo. High-density silicon probes such as the Neuropixel Ultra [ye2024ultra] can record extracellular voltage traces across hundreds of closely spaced electrodes. These signals, which reflect the transmembrane currents generated by nearby neurons, are traditionally used for spike detection and sorting—identifying when and where neurons fire spikes [pachitariu2016fast, rossant2016spike].
@@
While effective for tracking population activity, this approach treats spikes as discrete events and ignores the exquisite biophysical details of neurons [einevoll2012towards].
In this work, we revisit the extracellular signal, not just as a source of discrete spike times, but as a rich, spatially structured observation from which we can infer biophysical properties of individual neurons. This is made possible by three key developments: (1) high-density probes and 2D arrays that allow us to sample extracellular voltage at many sites around a cell to triangulate its position and morphology; (2) differentiable simulators like Jaxley [deistler2024differentiable] that enable fast, hardware-accelerated simulation of complex biophysical models and gradient-based parameter optimization; and (3) scalable state-space inference frameworks such as Dynamax [linderman2025dynamax], which allow efficient state inference and parameter estimation in high-dimensional, nonlinear dynamical systems.
We leverage these developments to provide a statistical approach for inferring the parameters of multi-compartment biophysical models from extracellular observations: We extend classical multi-compartment Hodgkin–Huxley models to include a biophysical model of extracellular potentials. We then use an Extended Kalman Filter (EKF) [sarkka2023bayesian] to infer latent neuronal states—such as membrane voltages and gating variables—from observed extracellular signals and we improve the computational efficiency with diagonal and block-diagonal approximations to the filtering covariance. To estimate model parameters, including ion channel densities and the neuron’s spatial position relative to the probe, we maximize the marginal likelihood using gradient-based optimization.
@@ 2.1 Modeling Action Potentials with Hodgkin–Huxley Dynamics
where $I_{\text{ext}}^{(i)}$ is the external current applied to compartment $i$, and $g^{(i,j)}$ is the axial conductance coupling compartment $i$ to its neighbor $j$. The total transmembrane current in compartment $i$ is denoted by $I_{m}^{(i)}$ and is the sum of the ionic current $I_{\text{ion}}^{(i)}=S^{(i)}\sum_{c\in\mathcal{C}}g_{c}^{(i)}a_{c}(\lambda_{c}^{(i)})(V^{(i)}-E_{c})$} and the capacitive current $I_{\text{cap}}^{(i)}=S^{(i)}C_{m}\frac{\mathrm{d}V^{(i)}}{\mathrm{d}t}$ [gold2007using]. For instance, in the original single compartment HH model with sodium, potassium, and leak channels, the total current is,
$I_{m}=S\left(g_{\text{Na}}\,m^{3}h(V-E_{\text{Na}})+{g}_{\text{K}}\,n^{4}(V-E_{\text{K}})+{g}_{\text{L}}(V-E_{\text{L}})+C_{m}\frac{\mathrm{d}V}{\mathrm{d}t}\right).$ (2)
Here, for example, the states of the sodium channel are $\lambda_{\mathrm{Na}}=(m,h)\in[0,1]^{2}$ and the activation function is $a_{\mathrm{Na}}(\lambda_{\mathrm{Na}})=m^{3}h$. The gating variables evolve according to first-order kinetics $\frac{\mathrm{d}\lambda_{c}}{\mathrm{d}t}=\alpha_{c}(V)(1-\lambda_{c})-\beta_{c}(V)\lambda_{c},$ where $\alpha_{c}(V)$ and $\beta_{c}(V)$ are voltage-dependent rate functions that control the opening and closing transitions of the ion channel. Specific forms for the $\alpha$ and $\beta$ functions are available in dayan2005theoretical, e.g. Importantly, note that by eq. (1), the ionic current and the capacitive current cancel out to give the following expression of the transmembrane current:
$I_{m}^{(i)}=I_{\text{ext}}^{(i)}+S^{(i)}\sum_{j\sim i}g^{(i,j)}(V^{(j)}-V^{(i)}).$ (3)
@@ 2.2 Modeling Extracellular Voltage Measurements
Extracellular voltages reflect the collective activity of nearby neurons. While the physics of extracellular recordings are well understood [gold2006origin, linden2014lfpy], our work is, to our knowledge, the first to leverage this framework for fitting multi-compartment HH models to extracellular measurements. Let $\Phi$ denote the extracellular potential measured by a fixed site on a recording probe. We treat each compartment as a point source of transmembrane current located at its center and denote $r_{i}$ the distance from the recording site to compartment $i$ of a neuron. It follows from Coulomb’s law that the measured extracellular voltage is,
@@
In words, the observed potential $\Phi$ is proportional to the sum of transmembrane currents from nearby compartments, weighted by the inverse of the distance to the compartment. Importantly, it follows from eq. (3) and (4) that the extracellular voltages at time $t$ are simply linear functions of the membrane voltages at time $t$.
@@ 3 State Inference and Parameter Estimation
We aim to infer the membrane potential and channel states and to estimate the biophysical and structural parameters of a neuron from noisy extracellular voltage traces. We can achieve both goals by casting the multi-compartment HH model as a state space model (SSM) with nonlinear dynamics. However, the nature and scale of these models present unique challenges, which we address below.
@@ 3.1 State Space Model Formulation
A state space model (SSM) is a latent variable model with latent states $\mathbf{z}_{1:T}$, observations $\mathbf{y}_{1:T}$, and parameters $\theta$. In a Markovian SSM, latent states evolve according to a dynamics distribution, $p(\mathbf{z}_{t}\mid\mathbf{z}_{t-1};\theta)$, and they give rise to observations via an emission distribution $p(\mathbf{y}_{t}\mid\mathbf{z}_{t};\theta)$. There are a host of well-established techniques for state inference and parameter estimation in such models [sarkka2023bayesian, linderman2025dynamax], depending on the nature of the dynamics and emission distributions.
In our case, the model parameters, $\theta$, may include the maximum conductances of various ion channels as well as parameters governing the neuron’s spatial configuration relative to the recording probe. The latent states consist of the voltages and channel states for each compartment, $\mathbf{z}_{t}=[V^{(1)},\dots,V^{(N)},\{\lambda_{c}^{(1)}\}_{c\in\mathcal{C}_{1}},\dots,\{\lambda_{c}^{(N)}\}_{c\in\mathcal{C}_{N}}]$. To model the state dynamics, we use an Euler discretization of the multi-compartment HH differential equations (see eq.˜1),
$\mathbf{z}_{t+1}$ $\sim\mathcal{N}(f_{\theta}(\mathbf{z}_{t},t),\,\Delta t\,\mathbf{\Sigma}_{\text{dyn}}),$ (5)
where $f_{\theta}(\mathbf{z}_{t},t)$ is the deterministic update function derived from the ODEs above, and $\mathbf{\Sigma}_{\text{dyn}}$ is a diagonal covariance matrix with entries from $\{\sigma^{2}_{v},~\sigma^{2}_{\text{state}}\}$. This Gaussian noise is intended to account for model uncertainty, such as missing ion channels or misspecified channel dynamics, which we will consider in the results section.
We assume that the only source of stochasticity in the emission process is measurement noise. Let $\mathbf{y}_{t}\in\mathbb{R}^{M}$ denote the observed extracellular potential at time $t$ across $M$ recording sites. It is modeled as,
$\mathbf{y}_{t}\sim\mathcal{N}(h_{\theta}(\mathbf{z}_{t}),\sigma^{2}_{\text{obs}}\mathbf{I}),$ (6)
where $h_{\theta}(\mathbf{z}_{t})$ is the predicted extracellular voltage from eq.˜4 and $\sigma^{2}_{\text{obs}}$ is the observation noise variance. Since $h_{\theta}(\mathbf{z}_{t})$ only depends on $\mathbf{z}_{t}$, the emission model is conditionally independent of previous states given the current latent state.
@@ 3.2 Scaling the Extended Kalman Filter to Large Biophysical Models
Since the dynamics are nonlinear, there is no closed-form solution for the filtering distribution $p_{\theta}(\mathbf{z}_{1:t}\mid\mathbf{y}_{1:t})$. To address this challenge, we use the extended Kalman filter (EKF), which linearizes the dynamics and emission functions via first-order Taylor expansions [sarkka2023bayesian]. This linear approximation enables recursive estimation of the latent states, and it also provides an estimate of the marginal likelihood $p_{\theta}(\mathbf{y}_{1:T})$ for parameter estimation.
@@
Biologically realistic neuron models often contain dozens or even hundreds of compartments, leading to latent state dimensions in the hundreds or thousands. The standard EKF maintains an estimate of the full state covariance matrix, and updating the filtering distribution requires solving a linear system with this matrix. Since this computation scales cubically with the number of states, it becomes computationally infeasible to apply standard EKF formulations to full-scale neuronal models. To overcome this challenge, we develop two scalable EKF variants for biophysical neuron model fitting.
Diagonal approximation
Our first scalable variant adopts a diagonal approximation of the posterior covariance, inspired by chang2022diagonal, while exploiting the sparsity inherent in multi-compartment HH dynamics. Specifically, we approximate the EKF filtering distribution $p_{t}(\mathbf{z})=\mathcal{N}(\mathbf{z};\boldsymbol{\mu}_{t},\boldsymbol{\Sigma}_{t|t})$ with a diagonal Gaussian $q_{t}(\mathbf{z})=\mathcal{N}(\mathbf{z};\tilde{\boldsymbol{\mu}}_{t},\tilde{\boldsymbol{\Sigma}}_{t|t})$ where $\tilde{\boldsymbol{\Sigma}}_{t|t}$ and $\tilde{\boldsymbol{\mu}}_{t}$ are chosen to minimize the divergence $\mathrm{KL}(q_{t}\,\|\,p_{t})$. As shown by chang2022diagonal, this minimization admits a closed-form solution: the optimal diagonal Gaussian approximation $q_{t}$ matches the marginal means and precisions (inverse variances) of $p_{t}$. This leads to efficient recursive updates that preserve the EKF structure while requiring only diagonal storage—reducing memory cost from $\mathcal{O}(d^{2})$ to $\mathcal{O}(d)$ for a $d$-dimensional state.
In practice, this approximation requires evaluating only the diagonal elements of two matrices: (i) $\mathbf{H}_{t}^{\top}\mathbf{R}_{t}^{-1}\mathbf{H}_{t}$, where $\mathbf{H}_{t}=\mathrm{Jac}(h_{\theta})(\boldsymbol{\mu}_{t|t-1})$ is the Jacobian of the emission model and $\mathbf{R}_{t}=\sigma^{2}_{\mathrm{obs}}\mathbf{I}$ is the observation noise covariance; and (ii) the predicted precision $\mathbf{\Sigma}_{t|t-1}^{-1}$, with
$\mathbf{\Sigma}_{t|t-1}=\mathbf{F}_{t}\mathbf{\Sigma}_{t-1|t-1}\mathbf{F}_{t}^{\top}+\mathbf{Q}_{t},$
@@
where $\mathbf{F}_{t}=\mathrm{Jac}(f_{\theta})(\boldsymbol{\mu}_{t|t-1})$ is the Jacobian of the dynamics and $\mathbf{Q}_{t}=\Delta t\,\boldsymbol{\Sigma}_{\mathrm{dyn}}$ is the dynamics covariance (cf. Appendix B).
The diagonal of $\mathbf{H}_{t}^{\top}\mathbf{R}_{t}^{-1}\mathbf{H}_{t}$ can be computed in linear time with respect to the number of states [chang2022diagonal]. Furthermore, because the measurement model maps the membrane voltages to extracellular voltages through a fixed linear operator (eq.˜3 and (4)), $\mathbf{H}_{t}^{\top}\mathbf{R}_{t}^{-1}\mathbf{H}_{t}$ is time-invariant, allowing its diagonal to be computed once and reused throughout filtering. The remaining challenge lies in efficiently computing $\mathrm{diag}(\mathbf{\Sigma}_{t|t-1}^{-1})$. To simplify calculations, we make the additional assumption that $\mathbf{\Sigma}_{t|t-1}$ is diagonal, reducing the problem to computing $\mathrm{diag}(\mathbf{\Sigma}_{t|t-1})$. Here, we exploit the sparsity pattern of $\mathbf{F}_{t}$, which follows directly from the local coupling structure of the HH equations. While voltage updates may depend on all other states due to the implicit Euler solver used in Jaxley [deistler2024differentiable], gating variables typically depend only on their own previous value and the voltage of their corresponding compartment. This sparsity enables linear-time computation of $\mathrm{diag}(\mathbf{\Sigma}_{t|t-1}^{-1})$. In practice, we compute $\mathbf{F}_{t}$ efficiently using the SparseJac library [sparsejac2023], which constructs Jacobians directly from sparsity patterns without instantiating the full dense matrix.
When is the diagonal approximation reasonable? We assume that the initial filtered covariance matrix $\mathbf{\Sigma}_{0}$ is diagonal. The EKF covariance update is:
$\boldsymbol{\Sigma}_{t|t}=\left(\boldsymbol{\Sigma}_{t|t-1}^{-1}+\mathbf{H}_{t}^{\top}\mathbf{R}_{t}^{-1}\mathbf{H}_{t}\right)^{-1}.$
The procedure we described is equivalent to performing this update step with both $\boldsymbol{\Sigma}_{t|t-1}$ and $\mathbf{H}_{t}^{\top}\mathbf{R}_{t}^{-1}\mathbf{H}_{t}$ diagonal. Is this valid?
In Appendix B, we show that for well-behaved HH kinetics—smooth transition rates, small integration steps, and not too frequent spiking—the dynamics Jacobian is dominated by its diagonal. Under these conditions, a diagonal prior covariance and dynamics covariance yield an approximately diagonal predicted covariance $\boldsymbol{\Sigma}_{t|t-1}$.
The tighter constraint arises from the emission model. The time-invariance matrix $\mathbf{H}_{t}^{\top}\mathbf{R}_{t}^{-1}\mathbf{H}_{t}$ reflects how strongly electrode signals overlap across compartments. When a neuron is discretized into many compartments and the electrodes are densely packed around it—as in Neuropixels or similar high-density probes—its off-diagonal entries can become large, and approximating it by a diagonal can severely degrade performance (cf. Appendix B for a complete derivation). Thus, the diagonal EKF appears best suited for electrode configurations in which each sensor predominantly captures activity from only a few nearby compartments.
Block-diagonal approximation
The diagonal EKF procedure described above assumes $\mathbf{H}_{t}^{\top}\mathbf{R}_{t}^{-1}\mathbf{H}_{t}$ to be diagonal. When this assumption fails, the true update step no longer preserves diagonality even for well-behaved HH models; enforcing it by truncation discards key aspects of the geometry of the problem and harms accuracy.
To avoid this loss, we introduce a more structured approximation that remains consistent under EKF updates. Since the emission model depends only on voltages, $\mathbf{H}_{t}^{\top}\mathbf{R}_{t}^{-1}\mathbf{H}_{t}$ is block-diagonal, with a single nonzero block corresponding to the voltage dimensions. This motivates a block-diagonal covariance structure, where the EKF predicted and filtered covariances consist of a dense voltage block and diagonal blocks for all remaining (gating) states. Concretely, let the first $K$ dimensions of the latent state vector correspond to compartment voltages and the remaining ones to gating variables. At each time step $t$, we assume that both the predicted and filtered covariance matrices take the form
$\boldsymbol{\Sigma}_{t|\,\star}=\begin{bmatrix}\boldsymbol{\Sigma}^{vv}_{t|\,\star}&0\\[3.0pt] 0&\operatorname{diag}\!\big(\boldsymbol{\sigma}^{2}_{t|\,\star}\big)\end{bmatrix},\qquad\star\in\{t{-}1,t\},$
where $\boldsymbol{\Sigma}^{vv}_{t|\,\star}\!\in\!\mathbb{R}^{K\times K}$ denotes the dense voltage–voltage covariance block, and $\boldsymbol{\sigma}^{2}_{t|\,\star}$ collects the gating variances. We refer to this family of matrices as the BD($K$) family (block-diagonal with $K$ compartments).
We can prove that if the dynamics Jacobian $\mathbf{F}_{t}$ is BD($K$), then the BD($K$) family is closed under the EKF prediction and update: if the initial covariance matrix $\boldsymbol{\Sigma}_{0}$ is BD($K$), then subsequent predicted and filtered covariances remain BD($K$) (cf. Proposition 1). As with the diagonal approximation, assuming a BD($K$) structure for the Jacobian is reasonable for well-behaved HH models.
Computationally, the EKF update step with BD($K$) covariances reduces to (i) a $K\times K$ solve for the voltage block—since the measurement Jacobian $\mathbf{H}_{t}$ has support only on voltages—and (ii) element-wise variance updates for the diagonal gating blocks (cf. proof of Proposition 1). The prediction step similarly preserves structure by propagating $\boldsymbol{\Sigma}^{vv}_{t-1|t-1}$ through the voltage–voltage sub-Jacobian and updating gating variances via their local Jacobians, yielding overall complexity $\mathcal{O}(K^{3}+d)$ instead of $\mathcal{O}(d^{3})$ for a $d$-dimensional state (cf. proof of Proposition 1).